OpenStudy (anonymous):

What is the standard form equation of the line passing through (3, –2) with a slope of three-fifths?

OpenStudy (anonymous):

with a slope of 3/5

OpenStudy (anonymous):

y- (-2) = (3/5) (x - 3) multiply each term by 5 5y + 10 = 3(x-3) 5y + 10 = 3x - 9 5y - 3x = -19 or u can write it as 3x - 5y = 19